OpenStudy (anonymous):

How many molecules are there in 24 grams of FeF3?

OpenStudy (anonymous):

\[FeF _{3}\]

OpenStudy (nechirwan):

n=wt/mwt n=24/112.839 0.213 mole No,of =No of mole *Avogadro's number =0.213*6.0221415 × 10>23 =1.2827 x 10>23